### v1

# OmniVLM: A Token-Compressed, Sub-Billion-Parameter Vision-Language Model for Efficient On-Device Inference

Abstract
We present OmniVLM, a sub-billion-parameter vision-language model for efficient on-device inference. OmniVLM introduces a token compression mechanism that reduces visual token sequence length from 729 to 81 tokens, significantly reducing computational overhead while preserving visual-semantic fidelity. Through a multi-stage training pipeline of pretraining, supervised fine-tuning, and minimal-edit Direct Preference Optimization (DPO), OmniVLM matches the performance of larger models. On multiple benchmarks including ScienceQA, POPE, and MMMU, OmniVLM outperforms existing baselines like nanoLLAVA within a 968M-parameter footprint. Empirical results on the same laptop demonstrate 9.1x faster time-to-first-token (0.75s vs 6.82s) and 1.5x higher decoding speed (29.41 vs 19.20 tokens/s) compared to nanoLLAVA, enabling efficient deployment on edge devices. The model weights can be accessed on huggingface: https://huggingface.co/NexaAIDev/OmniVLM-968M, and the inference examples can be find in Appendix B.
22footnotetext: Email: alexchen@nexa.ai
1 Introduction
Vision-language models (VLMs) are vital to on-device artificial intelligence applications, encompassing tasks from basic image captioning and optical character recognition to advanced capabilities such as visual question answering, video comprehension, and user interface analysis. Deploying these models necessitates the effective integration of visual and textual processing while meeting the strict computational and memory limitations present in edge devices such as smartphones, laptops, and embedded systems. Industry applications of VLMs are abundant in recent years, MultiON uses VLM for result supervision in their LLM agent Q [MultiON team, 2024]. Meta incorporated VLM in the newly released Orion AR glass [Meta, 2024b]. Apple developed the Ferret UI [Li et al., 2024, You et al., 2025] based on multimodal LLMs. Tencent uses multimodal LLM for agent on smartphones, the AppAgent [Zhang et al., 2023]. The Qwen team also recently developed Qwen2-VL as general-purpose VLM agent [Wang et al., 2024b].
The productionization of VLMs for edge deployment presents three primary challenges. First, visual input tokenization introduces substantial computational overhead. According to OpenAI’s analysis [Community, 2023], processing a single $1024\times 1024$ pixel image necessitates 765 tokens, comprising 170 tokens per $512\times 512$ tile plus an additional 85 tokens for base processing overhead. Second, power consumption poses a critical constraint for deployment on energy-limited devices. Recent empirical studies from Meta reveal that a 7B-parameter model consumes approximately 0.7J per token [Liu et al., 2024]. Consequently, processing a $1024\times 1024$ pixel image (765 tokens) requires approximately 536 J—consuming over 1% of an iPhone’s battery capacity (50 kJ)—before accounting for text processing. Third, existing VLMs with parameters below 2B demonstrate limited visual comprehension capabilities: the 1B-parameter nanoLLAVA achieves only 28.6% accuracy on the MMMU benchmark [Nguyen, 2024], significantly underperforming compared to OpenAI’s O1 model at 92.3% [OpenAI, 2024]. These challenges underscore the complexity of optimizing the trade-off between computational efficiency, energy consumption, and model performance for practical edge deployments.
To address these limitations, we present OmniVLM, a sub-billion parameter (968M) multimodal model specifically optimized for edge device deployment. Our key contributions include: (1) a novel token compression mechanism that reduces image token requirements by 9x while maintaining visual fidelity, and (2) an enhanced output quality framework utilizing Direct Preference Optimization (DPO) with minimal parameter updates. Through comprehensive evaluation on established benchmarks including MM-VET and ScienceQA, we demonstrate that OmniVLM achieves competitive performance while maintaining deployment feasibility on resource-constrained devices. Our research advances the state-of-the-art in efficient vision-language modeling, enabling practical deployment of multimodal capabilities across a broad range of edge computing scenarios.
2 Related Works
2.1 Vision-Language Models: Evolution and Architectures
Building an instruction following agent has been an active field of study in computer vision. Vision language models have achieved impressive performance, since the introduction of LLaVA  [Liu et al., 2023b, a], a large multimodal model that combines a vision encoder and a language model to create an end-to-end trained system for general-purpose visual and language understanding. Their work generated multimodal instruction-following data using GPT-4, thereby addressing the scarcity of such data. Two evaluation benchmarks were introduced with diverse and challenging tasks to assess the model’s capabilities. Additionally, when fine-tuned on ScienceQA [Lu et al., 2022], a multimodal reasoning dataset, LLaVA in conjunction with GPT-4 achieves state-of-the-art accuracy. The research on vision language models has been further pushed by efforts in both academia and industries. Recent progresses include PaliGemma [Beyer et al., 2024] built on top of Gemma-2B [Team et al., 2024], Qwen2-VL [Bai et al., 2023, Wang et al., 2024a] built on top of QwenLM [Yang et al., 2024], Llama3.2-Vision built on top of Llama [Meta, 2024a, Dubey et al., 2024] and InternVL [Chen et al., 2024b] built from vision transformer [Dosovitskiy, 2020] with language models like QLLaMA [Cui et al., 2023] and Vicuna [Zheng et al., 2023] as middleware.
2.2 Edge Deployment and Inference Optimization
Due to constraints on memory and computing capability, deploying language models on edge devices such as smartphones and PC has been a challenge. Techniques such as quantization, pruning, and KV cache are widely used. The GGML library [Gerganov, 2023-2024a] provides necessary functions for fast inference of large language models and the gguf format for model weights, based on which abundant high-level inference code have been developed, such as llama.cpp [Gerganov, 2023-2024b] and ctransformers [Marella, 2023-2024]. Applications that are designed particularly for local runs have emerged in recent years. Notable softwares include Ollama [Ollama team, 2023-2024], MLC [MLC team, 2023-2024] and etc.
3 Methodology
3.1 Model Architecture Overview
The OmniVLM architecture expands on the standard LLaVA architecture [Liu et al., 2023b, a]. As depicted in Figure 1, the vision encoder converts images into embeddings, which are then adjusted by the projection layer to integrate seamlessly with the language model. Compared to the standard LLaVA design, our enhancements reduce computational overhead while maintaining performance.
  • Base Language Model: The Qwen2.5-0.5B-Instruct model processes text inputs with strong contextual understanding.
  • Vision Encoder: Google’s SigLIP-400M generates high-quality image embeddings at a resolution of $384\times 384$ with a $14\times 14$ patch size.
  • Projection Layer: A Multi-Layer Perceptron (MLP) aligns image embeddings to the token space of the base model. Our novel projection layer reduces the number of image tokens by 9×9$\times$9 ×, addressing the inefficiencies in the standard LLaVA architecture.
Figure 1: OmniVLM model architecture
3.2 Image Token Compression
[…]
To optimize the compression methodology, we conducted a comparative analysis of three distinct dimensional transformation techniques: reshaping and two convolution-based approaches. The first convolution variant implements a 1D architecture with kernel size $k=9$ and stride $s=9$ (no padding), while the second employs a 2D configuration with kernel size $(9,1)$ and corresponding stride $(9,1)$. Both convolution approaches were designed to reduce sequence dimensionality while preserving spatial information. However, our experimental results demonstrate that these convolution-based methods consistently produce higher validation losses/perplexity compared to the reshaping strategy.
[…]
3.3 Minimal-Edit DPO for Enhanced Response Quality
Direct Preference Optimization (DPO) has proven to be highly effective in Reinforcement Learning with Human Feedback (RLHF), significantly reducing hallucinations, improving response correctness, and enhancing model safety. We developed a minimal-edit DPO method, where a teacher model applies small, targeted edits to the base model’s outputs. These minimally altered pairs focus on precise improvements in response quality while preserving the original model’s behavior. This approach allows for more controlled optimization, ensuring the benefits of DPO without compromising the stability of the base model. An example DPO pair dataset is presented in the Appendix A.
3.4 Multi-Stage Training
We developed the OmniVLM model through a three-stage training pipeline:
  • Pretraining: The model is initialized with foundational visual-linguistic alignments using large-scale image-caption datasets. During this phase, only the projection layer is trained to efficiently learn the mapping between visual and textual domains. The primary objective was to establish strong visual-linguistic alignments through caption generation tasks, enabling the model to form basic associations between visual features and textual descriptions.
  • Supervised Fine-Tuning (SFT): The model is fine-tuned on structured datasets comprising image-based question-answer pairs. This stage improves contextual understanding and conversational coherence. This stage was designed to enhance the model’s contextual understanding and improve its response generation capabilities across diverse scenarios. We freeze the vision encoder and train the projector as well as the LLM part.
  • Direct Preference Optimization (DPO): The final stage employs minimal-edit DPO, where the teacher model corrects base outputs with minor edits. These corrections focus on accuracy-critical elements, forming chosen-rejected pairs for effective preference learning. This stage involved several key steps:
    • Deployment of a specialized teacher model to generate minimally edited corrections that preserve essential content while improving quality.
    • Creation of chosen-rejected pairs by systematically comparing original outputs with their corrected versions.
    • Application of DPO fine-tuning to align the model’s behavior with preferred outputs, effectively teaching it to generate higher-quality responses that match human preferences. In this stage, we keep vision encoder frozen and train both projector and LLM backbone.
4 Experiments
4.1 Training Dataset
We trained OmniVLM through a carefully structured three-stage process, leveraging diverse datasets to achieve robust multimodal capabilities:
  • Pretraining: In this foundational stage, we utilized large-scale image-caption pairs, primarily sourcing from the LLaVA pretraining dataset [Liu, 2023]. The training focused exclusively on optimizing the connector module, processing approximately 558K training samples.
  • Supervised Fine-Tuning (SFT): During this critical second stage, we expanded our focus to optimize the entire system architecture. The training incorporated 6M carefully curated samples from LLaVA dataset [Liu, 2023], UnimmChat dataset [Yirany, 2024] and internal dataset from Nexa AI.
  • Direct Preference Optimization (DPO): In the final stage, we leveraged image and prompt data from the RLAIF-V project [OpenBMB, 2024] with reject and accepted pairs, and the rejected text results are generated from the model before DPO.
4.2 Compression Ratio Analysis
We conducted a systematic evaluation of different token compression ratios to determine the optimal balance between computational efficiency and model performance. Starting from the original 729 tokens generated by the SigLip-384 module, we experimented with four compression targets: 729 (baseline), 243 (3× reduction), 81 (9× reduction), and 9 (81× reduction).
The validation loss curves for different compression ratios are shown in Figure 2. Our experiments revealed that the 81-token configuration (9× compression) achieves the best performance among all tested ratios. This finding aligns with the architectural constraints of our small-scale language model backbone, where excessive token sequences can degrade performance due to attention mechanism limitations and context window constraints.
Figure 2: Validation loss curves across different image token compression ratios, evaluated on a test dataset comprising around 500K text-image pairs. The comparison demonstrates the effect of token reduction from the baseline (729 tokens) to various compression levels (243, 81, and 9 tokens).
Our analysis showed that the 729-token baseline exhibited higher validation losses, likely due to the computational burden on the attention mechanisms. While the 243-token configuration showed improved performance, it still retained unnecessary redundancy in the visual representation. The 81-token setting achieved an optimal balance, preserving essential visual information while minimizing computational overhead. In contrast, the 9-token configuration proved too aggressive, resulting in significant information loss and degraded performance across our evaluation metrics.
These results demonstrate that moderate token compression not only reduces computational requirements but can actually enhance model performance by preventing attention mechanism saturation in resource-constrained architectures. The 9× compression ratio (81 tokens) emerged as the sweet spot for our model, offering substantial efficiency gains without compromising visual understanding capabilities. This finding suggests that for compact vision-language models, aggressive token compression up to a certain threshold can be beneficial for both computational efficiency and model performance.
4.3 Quality Benchmark
Below, we present how OmniVLM outperforms nanoLLAVA in Table 1, the previous sub-billion vision-language model in open-source community. OmniVLM consistently achieves higher scores than nanoLLAVA, showcasing its enhanced reasoning capabilities, multimodal comprehension, and generalization across diverse tasks, as shown in Figure 3.
Specifically, on the ScienceQA benchmark, OmniVLM achieves a significant improvement, with an evaluation and test score of 71.0, compared to nanoLLAVA’s 59.0. This underscores OmniVLM’s ability to better understand and reason about complex science questions that require multimodal input processing. On POPE, OmniVLM exhibits a remarkable gain, achieving 93.3 compared to nanoLLAVA’s 84.1, reflecting its superior performance in precision-oriented tasks. For MM-VET, which evaluates multimodal visual and textual understanding, OmniVLM scores 30.9, outperforming nanoLLAVA by 7.0 points. Similarly, on the MMMU test and evaluation benchmarks, OmniVLM surpasses nanoLLAVA by 13.5 and 9.6 points, respectively, demonstrating its strong multimodal reasoning and contextual integration.
Benchmark OmniVLM nanoLLAVA
ScienceQA (Eval) 71.0 59.0
ScienceQA (Test) 71.0 59.0
POPE 93.3 84.1
MM-VET 30.9 23.9
MMMU (Test) 42.1 28.6
MMMU (Eval) 40.0 30.4
Table 1: Performance comparison of OmniVLM and nanoLLAVA on various benchmarks.
Figure 3: OmniVLM model benchmark.
4.4 Performance on Edge Devices
We assessed the deployment efficiency of the model on edge devices by evaluating key metrics during local CPU and GPU inference. These metrics include Time-to-First-Token (TTFT)—the latency to generate the first token—and Decoding Speed—tokens generated per second.
4.4.1 Laptop Performance: AMD Ryzen AI
Benchmarks were conducted on the ASUS Zenbook S 16 (UM5606), equipped with the AMD Ryzen™ AI processor, under the following configuration:
  • CPU: AMD Ryzen™ 9 HX 370 (12 cores, 24 threads, 2.0–5.1 GHz, 36 MB cache).
  • GPU: AMD Radeon™ 890M.
  • Memory: 24 GB RAM, 1 TB SSD.
Table 2: Laptop Performance Metrics
Metric OmniVLM nanoLLAVA
TTFT (s) 0.75 6.82
Decoding Speed (tokens/s) 29.41 19.20
Performance analysis reveals significant differences between OmniVLM and nanoLLAVA models. OmniVLM demonstrates superior performance across both measured metrics. The Time To First Token (TTFT) for OmniVLM is notably faster at 0.75 seconds compared to nanoLLAVA’s 6.82 seconds, representing approximately a 9x improvement in initial response time. In terms of decoding speed, OmniVLM processes 29.41 tokens per second, while nanoLLAVA achieves 19.20 tokens per second, indicating approximately a 53% higher throughput for OmniVLM. These metrics suggest that OmniVLM offers more efficient real-time performance on the tested hardware configuration, particularly in scenarios where rapid response times are crucial.
4.4.2 Mobile Performance: SAMSUNG S22
The model’s performance was also evaluated on the Google Pixel 6 with the following specifications:
  • CPU: Octa-core (1x3.00 GHz Cortex-X2 & 3x2.50 GHz Cortex-A710 & 4x1.80 GHz Cortex-A510).
  • GPU: Adreno 730.
  • Memory: 8 GB RAM, 128 GB storage.
Table 3: Mobile Performance Metrics
Metric OmniVLM nanoLLAVA
TTFT (s) 7.48 60.23
Decoding Speed (tokens/s) 31.88 24.33
The results in Table 3 highlight the significant efficiency of OmniVLM over nanoLLAVA in both TTFT and decoding speed, demonstrating its suitability for latency-sensitive edge applications. On mobile devices, OmniVLM achieves a TTFT of 7.48 seconds, which is approximately 8 times faster than nanoLLAVA’s 60.23 seconds. This substantial reduction in initial response time is particularly crucial for mobile applications where user experience depends heavily on rapid system responsiveness. In terms of decoding speed, OmniVLM processes 31.88 tokens per second compared to nanoLLAVA’s 24.33 tokens per second, representing a 31% improvement in processing throughput. This enhanced decoding efficiency translates to smoother real-time interactions and better overall performance in resource-constrained mobile environments. The performance gains are particularly noteworthy considering the computational limitations typically associated with mobile hardware, suggesting that OmniVLM’s architecture is well-optimized for edge deployment scenarios.
5 Future Work and Conclusion
In this paper, we presented OmniVLM, a compact and efficient sub-billion parameter vision-language model designed for resource-constrained edge devices. By leveraging innovative techniques such as advanced image token compression, minimal-edit Direct Preference Optimization (DPO), and a multi-stage training pipeline, OmniVLM achieves state-of-the-art performance while drastically reducing computational and energy overhead. Our benchmarks validate OmniVLM’s ability to outperform existing models like nanoLLAVA across linguistic and visual understanding tasks, while maintaining practicality for deployment on edge hardware, including laptops and mobile devices.
Looking ahead, we aim to extend OmniVLM’s capabilities by optimizing its inference on NPUs (Neural Processing Units) to unlock greater efficiency and scalability across various edge platforms. This includes collaboration with hardware partners to further reduce latency, power consumption, and memory requirements. Our ongoing research seeks to bridge the gap between high-performance multimodal AI and real-world applications, setting a foundation for the next generation of on-device intelligent systems.
Appendix A DPO Dataset Preparation
Our dataset preparation process leverages GPT-4V for generating synthetic training pairs with minimal edit distance. We applied the approach to generate a series of datasets for several stages of DPO training. For instance, we used GPT-4V-generated image caption data in ALLaVA dataset [Chen et al., 2024a] and generated DPO dataset through minimal edits on the captions. Table 4 lists two examples of the generated data.
prompt Please provide a detailed narrative of the image. Please illustrate the image through your words.
image [Uncaptioned image] [Uncaptioned image]
chosen The image displays a schematic representation of a quilted grid pattern with blue diagonal lines on a white background, symbolizing fabric, likely silk, due to the context of the question. Each grid section is separated by thin white lines, and a red-orange double border frames the entire grid. The top right corner has text indicating a "1" x 1" space", suggesting measurements for the grid. There’s no physical garment or ironing tools visible, which implies that the image is a conceptual or instructional illustration rather than a photograph of the actual task. The image is a section of an old map, presented in sepia tones indicative of aged parchment. It features handwritten labels, stylized iconography for trees, and several delineated plots of land, typical of historical land grant maps. Notable names are marked alongside illustrations of houses, perhaps indicating property ownership. A prominent river snakes through the center with tributary streams, and the scale suggests a close-up of a rural area. The absence of grid lines or latitude and longitude marks suggests it predates modern cartographic conventions.
rejected The image displays a diagram or illustration of a grid of blue and white squares, arranged in a pattern that resembles a solar panel array. The squares are arranged in a 3x3 grid, with each square containing a blue color and white lines. The blue squares are likely representing the photovoltaic cells, while the white lines may indicate the conductive pathways or the edges of the solar panels. The image also contains a red border around the grid, and there is a note indicating a "1" x 1" space, suggesting a specific measurement or spacing between the cells. The overall image appears to be a step-by-step guide or instructional material, possibly related to solar panel installation or maintenance. The image is a map, depicting a geographic area with various geographic features and locations. The map is detailed, showing topographical elements such as rivers, mountains, and landforms. There are several labeled areas, including "Griffin," "Cottwood," and "M. McQuarrie." The map is colored in shades of green, brown, and beige, with the green representing land and the brown and beige indicating water bodies. The map is not a satellite view, but rather a more detailed and close-up view of the area. There are no people or animals visible in the image.
Table 4: examples of DPO dataset generated from allava_vflan image caption data with minimal edit distance.
Appendix B Image Inference Examples
Based on community feedback, we continuously refined our model across five key dimensions:
  • Artistic Description Enhancement: Improved recognition and description of artistic styles, techniques, and compositional elements
  • Complex Scene Analysis: Advanced capabilities in understanding multi-object relationships and contextual interactions
  • Anime/Manga Recognition: Specialized improvements in identifying anime-specific visual elements and character traits
  • Detail Perception: Enhanced accuracy in color detection, texture analysis, and fine-grained visual feature recognition
  • Knowledge Integration: Expanded contextual understanding through integration of broader world knowledge
The following examples demonstrate OmniVLM’s capabilities across these dimensions.
Figure 4: Example of art description.
Figure 5: Example of complex scene analysis.
Figure 6: Example of anime recognition.
Figure 7: Example of detail perception.
Figure 8: Example of world knowledge.

### v2

@@
 The research on vision language models has been further pushed by efforts in both academia and industries.
-Recent progresses include PaliGemma [Beyer et al., 2024] built on top of Gemma-2B [Team et al., 2024], Qwen2-VL [Bai et al., 2023, Wang et al., 2024a] built on top of QwenLM [Yang et al., 2024], Llama3.2-Vision built on top of Llama [Meta, 2024a, Dubey et al., 2024] and InternVL [Chen et al., 2024b] built from vision transformer [Dosovitskiy, 2020] with language models like QLLaMA [Cui et al., 2023] and Vicuna [Zheng et al., 2023] as middleware.
+Recent progresses include PaliGemma [Beyer et al., 2024] built on top of Gemma-2B [Team et al., 2024], Qwen2-VL [Bai et al., 2023, Wang et al., 2024a] built on top of QwenLM [Yang et al., 2024], Llama3.2-Vision built on top of Llama [Meta, 2024a, Dubey et al., 2024] and InternVL [Chen et al., 2024c] built from vision transformer [Dosovitskiy, 2020] with language models like QLLaMA [Cui et al., 2023] and Vicuna [Zheng et al., 2023] as middleware.
 2.2 Edge Deployment and Inference Optimization
@@
 The GGML library [Gerganov, 2023-2024a] provides necessary functions for fast inference of large language models and the gguf format for model weights, based on which abundant high-level inference code have been developed, such as llama.cpp [Gerganov, 2023-2024b] and ctransformers [Marella, 2023-2024].
-Applications that are designed particularly for local runs have emerged in recent years.
+Applications [Chen and Li, 2024, Chen et al., 2024b] that are designed particularly for local runs have emerged in recent years.
 Notable softwares include Ollama [Ollama team, 2023-2024], MLC [MLC team, 2023-2024] and etc.